40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15491     Entry time: Fri Jul 17 00:18:13 2020
Author: gautam 
Type: Update 
Category: General 
Subject: Locking updat 
  1. I found that an EPICS channel wasn't reset to the correct value by burtrestore after the FE bootfest yesterday.
    • This cost me the whole of last night, found it finally tonight. 
    • I'll try and modify the locking scripts to better capture such errors, but ideally, we should just use Guardian or something since it's made for this purpose already.
    • Anyways, tonight I was able to re-acquire the PRFPMI lock in a completely scripted way.
  2. Locking CARM on POX remains out of reach.
    • I think this has to do with the fact that the zero-crossing of the CARM and REFL error signals are dependent on the 3f PRCL/MICH error point offsets.
    • So even if the DC gain is right, the fact that we use POX for the digital AO path and REFL for the analog AO path is leading to some conflict I think.
    • Ran out of energy tonight, I'll try again tomorrow.

The DQ channels of the ETM coils were active tonight, so I'll make the coil driver actuation budget over the next couple of days.

ELOG V3.1.3-